OpenStudy (anonymous):

A diameter of a circle has endpoints at (4, 6) and the origin. Which point is also on the circle? A. (–1, 2) B. (6, 3) C. (5, 0) D. (–1, 1) E. NOTA Please explain.

OpenStudy (campbell_st):

find the midpoint x = (4 + 0)/2 y = (6 + 0)/2 that will give the centre

OpenStudy (campbell_st):

then the equation is (x - h)^2 + (y-k)^2 = r^2 in this case r^2 = 52 (h, k) is the centre... sub in each point... to test

OpenStudy (anonymous):

answer is D (-1,1) please can you show more steps

OpenStudy (anonymous):

I didn't get it

Directrix (directrix):